### v2

@@ Abstract
We examine core concepts in the classical model of shapley1974cores under full preferences. Among the standard notions, the strong core may be empty, whereas the weak core, though always nonempty, can be overly large and include unreasonable allocations. Our main findings are: (1) The exclusion core of balbuzanov2019endowments—a recent concept shown to outperform standard cores in complex environments under strict preferences—can also be empty. We establish a necessary and sufficient condition for its nonemptiness, showing that it is more often nonempty than the strong core. (2) We introduce two new core concepts, built on the exclusion core and the strong core respectively, by refining the assumptions on how indifferent agents may block. Both are nonempty and Pareto efficient, and coincide with the strong core whenever the latter is nonempty. (3) These core concepts are ordered by set inclusion, with the strong core as the smallest and the weak core as the largest.
@@ 1 Introduction
Understanding market design models from a cooperative perspective has been a tradition in the field, dating back to the seminal contributions of gale1962college and shapley1974cores. gale1962college introduce the concept of stable matching in the two-sided marriage model and establish its existence. shapley1974cores introduce the one-sided housing market model and prove the nonemptiness of the core, latter referred to as the weak core. The weak core is based on strong domination and may include Pareto inefficient allocations. In contrast, the strong core, defined via weak domination, has been favored in the field because it guarantees Pareto efficiency, a key desideratum in many market design environments. However, the strong core may be empty when agents have weak preferences. Since the housing market model has been extensively studied and played a foundational role in modeling various market design problems, this paper investigates these issues and seeks appropriate core concepts that ensure both nonemptiness and Pareto efficiency in the model.
@@
The weak core and the strong core differ in their assumptions on the blocking behavior of indifferent agents. In the weak core, an agent joins a blocking coalition only if he strictly benefits from the deviation, whereas the strong core permits indifferent agents to participate. This difference is inconsequential in transferable utility environments, since transfers within a blocking coalition can make all members strictly better off. However, it becomes crucial in market design settings where transfers are typically prohibited. Allowing indifferent agents to join blocking coalitions is essential for achieving Pareto efficiency, yet doing so may render the strong core empty. Because Pareto efficiency is central in market design, the literature often assumes strict preferences, under which the strong core is nonempty and includes a unique element in the housing market model, and can be computed via the top trading cycle (TTC) mechanism. Nevertheless, weak preferences are pervasive in practice and have been considered since the original work of shapley1974cores. For example, in kidney exchange, it is standard to assume that patients have weak preferences over compatible kidneys. In allocation problems with multi-unit objects, agents are naturally indifferent among identical copies of the same object. Additional arguments for studying weak preferences can be found in bogomolnaia2004random, bogomolnaia2005strategy, and erdil2017two, among others. We present the housing market model in Section 2 and define standard core concepts in Section 3.1.
@@
We approach the problem by examining the behavioral foundations of coalition formation under standard core concepts. The weak core adopts the standard economics assumption that agents are self-interested. In contrast, the strong core assumes that agents are “altruistic”: they may join a blocking coalition to benefit others as long as they are not harmed. This interpretation, however, is controversial, since indifferent agents may help someones in the coalition while potentially harming outside agents. We note that the notion of exclusion rights, recently introduced by balbuzanov2019endowments (hereafter BK), offers a new and more compelling interpretation. In the housing market model, a group of agents directly controls their endowments and holds the right to evict others who occupy their endowments. By threatening eviction, the group gains indirect control over those agents’ endowments. Leveraging these indirect control rights, the group can expand its control over more objects. Under strict preferences, in any blocking coalition that does not include redundant indifferent agents, [^1] In a blocking coalition, a group of indifferent agents is redundant if they allocate their endowments among themselves and thus contribute nothing to the others in the coalition. the object received by each indifferent agent must be directly or indirectly controlled by the strictly better-off agents. Thus, indifferent agents can be viewed as being compelled to join the coalition under the threat imposed by the better-off agents. This insight motivates us to build on exclusion rights to understand coalition formation under weak preferences.
@@
We first examine the exclusion core, a core concept based on exclusion rights, which is originally defined by BK for their general model that subsumes the housing market model as a special case. The exclusion core consists of allocations in which no coalition can strictly benefit all of its members by evicting others from the objects they control. BK show that, under strict preferences, the exclusion core is more effective than the strong core in eliminating unintuitive allocations in their model, and the two cores coincide in the housing market model. BK do not address weak preferences. Thus, we examine this concept as a potential new solution in the housing market model under weak preferences.
@@
Our examination of the exclusion core under weak preferences is new to the literature. We obtain the following results. First, it no longer coincides with the strong core and may also be empty. Second, we identify a necessary and sufficient condition for its nonemptiness (Proposition 1), which is strictly weaker than the counterpart for the strong core (quint2004houseswapping). So, the exclusion core is more often nonempty than the strong core. Moreover, whenever the strong core is nonempty, they coincide. Finally, the exclusion core fails to satisfy an intuitive property we call equivalence-closedness (Proposition 2), which requires that, for any two allocations that are indifferent to all agents, if one belongs to a solution, the other also belongs to the solution. This property is intuitive, as agents’ blocking decisions should depend only on their welfare. Both the strong core and the weak core satisfy this property. These results are presented in Section 3.2.
@@
The inadequacy of the exclusion core suggests us to reevaluate its definition under weak preferences, which leads to our revision of the concept in Section 4. Our first observation is that the notion of control rights needs to be revised under weak preferences. In BK’s definition, if an agent receives an object owned by a group, the group indirectly controls the agent’s endowment via threatening to evict him from that object. This threat is credible under strict preferences, since the agent cannot find other indifferent objects. Under weak preferences, however, since the agent may view several objects as indifferent, the threat becomes not credible if there exist other indifferent objects not owned by the group. We therefore redefine control rights as follows: a group directly controls their own endowments, and indirectly controls another agent’s endowment if and only if all objects that the agent views as indifferent to his assignment are controlled by the group.
@@
Our second observation is that the exercise of exclusion rights under weak preferences is more subtle than under strict preferences, which requires the inclusion of indifferent agents in blocking coalitions. This is in contrast to BK’s original definition, which precludes indifferent agents from joining blocking coalitions, to ensure the nonemptiness of the exclusion core in their model. In the housing market model, under weak preferences, subtle cases arise when a group of agents $I$ does not control another group $J$’s endowments but can still effectively evict others from $J$’s endowments. In these cases, all objects that $J$ views as indifferent to their current assignments but are not owned by $J$ are owned by $I$. In one case, if $I$ evicts $J$ from $I$’s endowments, $J$ is able to revert to their own endowments to maintain their welfare, which effectively allows $I$ to indirectly evict others from $J$’s endowments. In another case, if $I$ evicts $J$ from $I$’s endowments, $J$ is unable to maintain their welfare by reverting to their own endowments. That is, $J$ relies on $I$’s endowments to maintain their welfare. This dependence gives $I$ leverage to compel $J$ to join their coalition. Our new definition unifies these cases by requiring that, if a blocking coalition includes indifferent agents, then, for each indifferent agent, all indifferent objects not owned by their group must be exclusively owned by the group of strictly better-off agents in the coalition. Equivalently, all indifferent objects for each indifferent agent in the coalition are owned by the coalition.
@@
Our first solution concept, the rectified exclusion core, requires that if a coalition blocks an allocation via another one, then (1) at least one member of the coalition is strictly better off and no member is worse off, (2) every agent harmed by the blocking must be evicted from an object controlled by the coalition, and (3) if the coalition includes indifferent agents, all indifferent objects for indifferent agents must be owned by the coalition. This definition differs from the exclusion core in that we redefine controlled objects and allow indifferent agents to join a blocking coalition under condition (3). The rectified exclusion core is nonempty, Pareto efficient, and equivalence-closed (Theorem 1). It is a superset of the exclusion core and a subset of the weak core, and coincides with the strong core whenever the strong core is nonempty. To prove its nonemptiness, we propose a generalization of TTC to find its elements (Lemma 1). Our algorithm follows the procedures of the existing generalizations of TTC in the literature (alcalde2011exchange; jaramillo2012difference; aziz2012housing; saban2013house; plaxton2013simple; ahmad2021group). To preserve strategy-proofness of TTC, these algorithms need to carefully select pointing rules for agents. Since our purpose is only to find elements of our solution, we do not select a specific pointing rule, only requiring that at least one trading cycle is generated in each step. This implies that the outcomes of these existing algorithms all belong to the rectified exclusion core. Previously, the literature only shows that the outcomes of these algorithms belong to the weak core and are in the strong core when the strong core is nonempty.
@@
The notion of exclusion rights is a useful conceptual tool for understanding coalition formation under weak preferences, which forms the basis for our modification of the exclusion core. However, different from the exclusion core and our modification that view endowments as a distribution of exclusion rights, standard core concepts view endowments as something that agents must consume by themselves or exchange with others. So, they require that a blocking coalition must deviate by reallocating their endowments among themselves. By imposing this requirement while preserving the above condition for indifferent agents in a blocking coalition, we introduce another solution concept, the rectified strong core, in Section 5. It differs from the strong core in one critical aspect: while the strong core assumes that a single indifferent object suffices to justify the participation of an indifferent agent in a blocking coalition, the rectified strong core requires that all indifferent objects for the agent must be owned by the coalition. This distinction vanishes under strict preferences, but under weak preferences, it reflects opposing extreme assumptions on the blocking behavior of indifferent agents. The rectified strong core is nonempty, Pareto efficient, and equivalence-closed (Proposition 3). It is a superset of the rectified exclusion core and a subset of the weak core. It coincides with the strong core whenever the latter is nonempty.
Thus, we obtain the following relationships between the various core concepts:
@@
Except for the weak core, all others coincide with the strong core whenever the latter is nonempty. In particular, they coincide under strict preferences.
We further examine a special case of our model in Section 6, where agents’ indifferent preferences arise from the existence of multiple copies of objects. So, an agent views any two objects as indifferent if and only if all agents view the two objects as indifferent. In this special model, we show that the strong core may still be empty, whereas the exclusion core is nonempty and coincides with the set of outcomes of TTC in an artificial priority-based allocation model where each object ranks its owners by an arbitrary order. However, in this special model, the exclusion core is still not equivalence-closed. So, the rectified exclusion core can still be strictly larger than the exclusion core.
@@
We conclude the paper in Section 7 by discussing other potential solution concepts. First, wako2007nonexistence have shown that, in general, the von Neumann-Morgenstern (vNM) stable set, defined based on either strong domination or weak domination, does not exist in the housing market model under weak preferences. Second, while demuynck2019myopic have shown that their myopic stable set, defined based on weak domination, coincides with the strong core in the housing market model under strict preferences, we show that, under weak preferences, this solution may include Pareto inefficient allocations. Finally, although the bargaining set introduced by yilmaz2022stability is nonempty, Pareto efficient, and lies between the strong core and the weak core, we show that it can be strictly larger than the strong core when the latter is nonempty. So, it may include unintuitive allocations when obviously intuitive allocations exist.
Related literature.
@@
The housing market model has been extensively studied. Although shapley1974cores and roth1977weak, the two seminal papers on this model, both consider full preferences, follow-up papers have mostly focused on strict preferences; see afacan2024housing for a survey. Among the papers that have addressed weak preferences, some focus on generalizing TTC to maintain its Pareto efficiency and strategy-proofness properties that hold under strict preferences, and others are interested in the set of competitive allocations, because it is nonempty and lies between the strong core and the weak core. However, we do not think the set of competitive allocations is a desirable solution, since it often includes Pareto inefficient elements, can be strictly larger than the strong core when the latter is nonempty, and has anomalies that have been pointed out by roth1977weak. [^2] In some markets, all Pareto efficient allocations in the weak core are not competitive, all competitive allocations are not Pareto efficient, and there exist two allocations that are indifferent to all agents, but one is competitive yet the other is not.
@@
Specifically, wako1984note proves that a nonempty strong core may be a strict subset of the set of competitive allocations. wako1991some proves that the strong core coincides with the set of competitive allocations if and only if any two competitive allocations are indifferent to all agents. quint2004houseswapping provide a necessary and sufficient condition for the strong core to be nonempty, which implies that a nonempty strong core is essentially single-valued. biro2023shapley prove that the strong core satisfies the respect-improvement property if being nonempty, and the set of competitive allocations satisfies the property in a stochastic-dominance sense. [^3] The respect-improvement property requires that an agent must receive a weakly better assignment when his endowment becomes more desirable for others.
@@
Several papers have provided cooperative foundations for the set of competitive allocations. wako1999coalition shows that this set coincides with a modification of the strong core based on antisymmetric weak domination. [^4] Indifferent agents in a blocking coalition must receive the same object in the old and new allocations toda1997implementation shows that this set is the unique vNM stable set based on Wako’s antisymmetric weak domination. kawasaki2010farsighted and klaus2010farsighted respectively prove that this set is the unique vNM stable set based on a farsighted version of Wako’s antisymmetric weak domination, or based on a farsighted version of strong domination if no agent is indifferent between his endowment and any other object. In these results, additional conditions must be imposed on the domination relation or on agents’ preferences because a vNM-stable set generally does not exist (wako2007nonexistence). ehlers2004monotonic shows that the strong core is not Nash implementable and the minimal monotonic extension of the strong core is the set of competitive allocations. [^5] A monotonic extension of the strong core is a superset of the strong core that satisfies Maskin’s Monotonicty, which is necessary for Nash implementation.
@@
alcalde2011exchange and jaramillo2012difference independently propose the first generalizations of TTC to full preferences that preserve Pareto efficiency and strategy-proofness. aziz2012housing study the computational complexity of these algorithms, and plaxton2013simple proposes another algorithm with enhanced computational efficiency. saban2013house unify these generalizations by defining a family of algorithms and derive sufficient conditions on pointing rules to preserve strategy-proofness. ahmad2021group study weak group strategy-proofness of these algorithms. Our generalization of TTC to find elements of our proposed cores follow the common formats of these existing algorithms. But since we do not pursue strategy-proofness, our description shows more flexibility in selecting pointing rules.
@@
There is a literature studying standard cores in the housing market model with externalities (mumcu2007core; graziano2020shapley; hong2022core; dougan2011core; aslan2020competitive; klaus2023core). It shows that even the weak core can be empty under unrestricted strict preferences. So, it typically focuses on restricted preference domains on which the strong core or other solutions are nonempty.
Finally, following BK, a few papers have further studied the exclusion core. balbuzanov2019property generalize their concept from their original model to a more general one that incorporates production. ishida2025group study a special case of BK’s model, where agents are partitioned into groups and each group collectively owns a partitioned group of objects, and show that the strong core and the exclusion core have a nonempty intersection, though they do not include each other. zhangrefinedexclusioncore studies refinements of the exclusion core in BK’s model. All those papers assume strict preferences. In future research we may explore how to employ the approach in this paper to address weak preferences in those papers.

### v3

@@ Abstract
We examine core concepts in the classical model of Shapley and Scarf (1974) under full preferences. Among the standard concepts, the strong core may be empty, whereas the nonempty weak core may be overly large and contain inefficient elements. Our main findings are: (1) The exclusion core of Balbuzanov and Kotowski (2019)—a recent concept outperforming standard concepts in complex environments under strict preferences—can also be empty, yet it is more often nonempty than the strong core. (2) We introduce two new core concepts, respectively built on the exclusion core and the strong core. Both are nonempty and Pareto efficient, and coincide with the strong core whenever it is nonempty. (3) These core concepts are ordered by set inclusion, with the strong core as the smallest and the weak core as the largest.
@@ 1 Introduction
Understanding matching models from a cooperative perspective has been a long-standing tradition in the market design literature, dating back to the seminal contributions of Gale and Shapley (1962) and Shapley and Scarf (1974). Gale and Shapley (1962) introduce the two-sided matching model and the concept of stable matching. Shapley and Scarf (1974) introduce the one-sided assignment model (later referred to as the housing market model), which has since become the foundation of a wide range of allocation problems. In the model, unit-demand agents exchange their indivisible endowments based on ordinal preferences without involving money. The work of Shapley and Scarf (1974) has been well known for its introduction of the top trading cycles (TTC) mechanism, but their original purpose is to prove the nonemptiness of the core in the presence of indivisibility. The core concept they prove is the weak core, which consists of allocations in which no coalition of agents can strictly benefit from a reallocation of their endowments among themselves. The more demanding strong core consists of allocations in which no coalition weakly benefits from a reallocation of their endowments and at least one member strictly benefit. The weak core is nonempty, but it may be overly large and contain Pareto inefficient elements, whereas the strong core is Pareto efficient but may be empty when agents have weak preferences. Pareto efficiency is a key desideratum in many market design settings, and weak preferences are pervasive in practice and have been considered since the original work of Shapley and Scarf (1974). Compared to the extensive study of the TTC mechanism, the issues with the standard core concepts in the housing market model have received relatively limited attention. Given the foundational role of the model in the literature, this paper revisits these issues and seeks alternative core concepts that ensure both nonemptiness and Pareto efficiency when agents have unrestricted preferences.
@@
We approach the problem by examining the behavioral foundations of coalition formation under the standard core concepts. The weak core adopts the standard economics assumption that agents are self-interested: an agent joins a blocking coalition only if he strictly benefits. In contrast, the strong core assumes that agents are “altruistic” in the sense that they may join a blocking coalition to benefit others as long as they are not harmed. This difference is inconsequential in transferable utility settings, since transfers within a blocking coalition can make all members strictly better off. However, it becomes crucial in non-transferable utility settings. In the housing market model, allowing indifferent agents to join blocking coalitions is essential for guaranteeing Pareto efficiency, yet doing so may render the strong core empty. To solve this, the literature often restricts attention to strict preferences, under which the strong core is nonempty and contains a unique element, which can be found via TTC. Nevertheless, the strict preference assumption rules out many important applications that involve weak preferences. For example, in kidney exchange, it is standard to assume that patients are indifferent among compatible kidneys (Roth et al., 2005). In allocations of multi-unit objects, agents are naturally indifferent among identical copies of the same object. Additional arguments for studying weak preferences can be found in Bogomolnaia and Moulin (2004), Bogomolnaia et al. (2005), and Erdil and Ergin (2017), among many others.
@@
The altruism interpretation for the strong core is controversial, since indifferent agents may help someones in the blocking coalition while potentially harming outside agents. We note that the concept of exclusion rights recently introduced by Balbuzanov and Kotowski (2019) (hereafter BK) offers a more compelling interpretation. In the housing market model, a group of agents directly controls their endowments and holds the right to evict others who occupy their endowments. By leveraging this exclusion right, the group can gain indirect control of others’ endowments. Under strict preferences, in any blocking coalition in which no indifferent agents are redundant, [^1] In a blocking coalition, we view a group of indifferent agents as redundant if they allocate their endowments among themselves, so that they contribute nothing to the others in the coalition. the object received by each indifferent agent must be directly or indirectly controlled by the group of better-off agents. So, indifferent agents can be viewed as being pressured to join the coalition under the threat from better-off agents. This insight motivates us to build on exclusion rights to understand coalition formation in the housing market model under weak preferences.
@@
We first examine the exclusion core, a core concept originally proposed by BK for a general model that subsumes the housing market model as a special case. The exclusion core consists of allocations in which no coalition can strictly benefit all of its members by evicting others from its controlled objects. BK show that, under strict preferences, the exclusion core is more effective than the strong core in eliminating unintuitive allocations in the general model, and they coincide in the housing market model. BK do not study weak preferences. Given their results, it is natural to consider the exclusion core as a potential solution for the housing market model under weak preferences.
@@
We obtain the following results for the exclusion core under weak preferences. First, the exclusion core no longer coincides with the strong core and may also be empty. Second, we identify a necessary condition and a separate sufficient condition for an allocation to be in the exclusion core (Lemma 1). The necessary condition suggests that every allocation in the exclusion core must be an outcome of TTC after breaking preference ties. The sufficient condition implies that the exclusion core is a superset of the strong core. We further examine a special case of the housing market model in which agents’ indifferent preferences arise from the presence of multiple copies of objects. We prove that the exclusion core is nonempty and coincides with the set of outcomes of TTC in this setting, whereas the strong core may still be empty. So, the exclusion core is more often nonempty than the strong core. Finally, the exclusion core fails to satisfy an intuitive property called equivalence-closedness, which requires that, for any two allocations that are indifferent to all agents, if one belongs to a core concept, the other also belongs to that core. This property is natural, as agents’ blocking decisions should depend only on their welfare. The strong core and the weak core satisfy this property.
@@
The useful role of exclusion rights in understanding coalition formation and yet the inadequacy of the exclusion core motive us to revisit its definition under weak preferences. Our first observation is that the definition of indirect exclusion needs to be revised under weak preferences. In BK’s definition, if an agent receives an object owned by a different group, the group indirectly controls the agent’s endowment via threatening to evict him from that object. This threat is credible under strict preferences, since the agent cannot find other indifferent objects to maintain his welfare. Under weak preferences, however, the threat may become not credible because the agent may view other objects not owned by the group as equally good. So, in our new definition, a group indirectly controls another agent’s endowment if and only if all objects that the agent views as indifferent to his assignment are controlled by that group.
@@
Our second observation is that the execution of exclusion rights under weak preferences becomes more subtle than under strict preferences. It requires the inclusion of indifferent agents in blocking coalitions under certain conditions. This is in contrast to BK’s original definition, which precludes indifferent agents from joining blocking coalitions to ensure the exclusion core to be nonempty in their general model. In the housing market model under weak preferences, subtleties arise when a group of agents $I$ does not control another group $J$’s endowments according to our new definition but can still effectively evict others from $J$’s endowments. In one case, if $I$ evicts $J$ from $I$’s endowments, $J$ is able to maintain their welfare by reverting to their own endowments, which effectively allows $I$ to evict others from $J$’s endowments. In another case, if $I$ evicts $J$ from $I$’s endowments, $J$ is unable to maintain their welfare by reverting to their own endowments, though $I$ does not control all objects that $J$ views as indifferent to their assignments. So, $J$ relies on $I$’s endowments to maintain their welfare. This dependence gives $I$ leverage to compel $J$ to join their coalition. Our new definition unifies these cases by requiring that, if a blocking coalition includes indifferent agents, then, for each indifferent agent, all indifferent objects not owned by their group must be exclusively owned by the group of better-off agents in the coalition. Equivalently, all indifferent objects for each indifferent agent in the coalition must be owned by the coalition.
@@
Our first solution concept, the rectified exclusion core, requires that if a coalition blocks an allocation via another one, then (1) at least one coalition member is strictly better off and no member is worse off, (2) every agent harmed by the blocking must be evicted from an object controlled by the coalition, and (3) if the coalition includes indifferent agents, all indifferent objects for indifferent agents must be owned by the coalition. The rectified exclusion core is nonempty, Pareto efficient, and equivalence-closed (Theorem 1). It is a superset of the exclusion core and a subset of the weak core, and coincides with the strong core whenever the strong core is nonempty. To prove its nonemptiness, we utilize a generalization of TTC to find its elements (Lemma 2). Our algorithm follows the procedures of the existing generalizations of TTC in the literature (Alcalde-Unzu and Molis, 2011; Jaramillo and Manjunath, 2012; Aziz and De Keijzer, 2012; Saban and Sethuraman, 2013; Plaxton, 2013; Ahmad, 2021). To preserve Pareto efficiency and strategy-proofness, these existing algorithms need to carefully select pointing rules for agents. Since our purpose is only to find elements of our solution, we do not select a specific pointing rule, only requiring that at least one trading cycle is generated in each step. So, our result implies that the outcomes of these existing algorithms all belong to the rectified exclusion core. Existing results only show that the outcomes of these algorithms are in the weak core and are in the strong core when the strong core is nonempty.
@@
Different from the exclusion core that interprets endowments as a distribution of exclusion rights, standard core concepts view endowments as something that agents must consume by themselves or exchange with others. So, a coalition is feasible to block an allocation only if its members reallocate their endowments among themselves. By imposing this feasibility requirement while preserving the condition under the rectified exclusion core for indifferent agents to join a blocking coalition, we introduce another solution concept called the rectified strong core. It differs from the strong core in one aspect: under the strong core, a single indifferent object assigned to an agent suffices to justify his participation in a blocking coalition, whereas the rectified strong core requires that all indifferent objects for the agent must be owned by the coalition. This distinction vanishes under strict preferences, but under weak preferences, it reflects opposing extreme assumptions on the blocking behavior of indifferent agents. The rectified strong core is nonempty, Pareto efficient, and equivalence-closed (Theorem 2). It is a superset of the rectified exclusion core and a subset of the weak core and coincides with the strong core when it is nonempty. So, we obtain the following relationships between the various core concepts:
@@
Except for the weak core, all others coincide with the strong core whenever it is nonempty, and they coincide under strict preferences.
@@
We conclude the paper by discussing other potential solutions. First, as Wako et al. (2007) have shown, in general, the von Neumann-Morgenstern (vNM) stable set based on either strong domination or weak domination does not exist in the housing market model under weak preferences. Second, while Demuynck et al. (2019) have shown that the myopic stable set based on weak domination coincides with the strong core in the housing market model under strict preferences, we show that, under weak preferences, the myopic stable set may contain Pareto inefficient allocations. Moreover, both the vNM stable set and the myopic stable set are setwise solutions; whether an allocation belongs to the set depends on other allocations in the set. In contrast, the core concepts we discuss are pointwise solutions. Finally, the bargaining set introduced by Yılmaz and Yılmaz (2022) is nonempty, Pareto efficient, and lies between the strong core and the weak core, but we show that it can be strictly larger than the strong core when the latter is nonempty.
The rest of the paper is organized as follows. After discussing related literature, we define the housing market model along with the standard core concepts in Section 2. Section 3 presents our results for the exclusion core. Section 4 introduces the rectified exclusion core. Section 5 introduces the rectified strong core. Section 6 discuss other potential solutions. All proofs are presented in the Appendix.
Related literature
@@
The housing market model has been extensively studied in the literature; see Afacan et al. (2024) for a survey. While Shapley and Scarf (1974) and Roth and Postlewaite (1977), the two seminal papers on this model, both consider full preferences, the literature has mostly focused on strict preferences. Among the papers that address weak preferences, some generalize TTC to maintain its Pareto efficiency and strategy-proofness properties, and others are interested in the set of competitive allocations, which coincides with the set of TTC outcomes with arbitrary tie breaking. However, the set of competitive allocations is not a desirable solution, since it often contains Pareto inefficient elements, can be strictly larger than a nonempty strong core, and has anomalies pointed out by Roth and Postlewaite (1977): in some markets, all competitive allocations are not Pareto efficient, and there exist two allocations that are indifferent to all agents, but one is competitive yet the other is not.
@@
Specifically, Wako (1984) proves that a nonempty strong core may be a strict subset of the set of competitive allocations, and Wako (1991) proves that they coincide if and only if any two competitive allocations are indifferent to all agents. Quint and Wako (2004) provide a necessary and sufficient condition for the strong core to be nonempty, implying that a nonempty strong core is essentially single-valued. We utilize this result when studying the relationship between the strong core and our core concepts. Biró et al. (2023) prove that the strong core satisfies the respect-improvement property if being nonempty and the set of competitive allocations satisfies the property in a stochastic-dominance sense. [^2] The respect-improvement property requires that an agent must receive a weakly better assignment when his endowment becomes more desirable for others.
@@
Several papers have provided cooperative foundations for the set of competitive allocations. Wako (1999) shows that this set coincides with a modification of the strong core based on antisymmetric weak domination. [^3] Indifferent agents in a blocking coalition must receive the same object in the old and new allocations. Toda (1997) shows that this set is the unique vNM stable set based on Wako’s antisymmetric weak domination. Kawasaki (2010) and Klaus et al. (2010) respectively prove that this set is the unique vNM stable set based on a farsighted version of Wako’s antisymmetric weak domination, or based on a farsighted version of strong domination if no agent is indifferent between his endowment and any other object. In these results, additional conditions must be imposed on the domination relation or on agents’ preferences, because a vNM-stable set generally does not exist (Wako et al., 2007). Ehlers (2004) shows that the strong core is not Nash implementable and the minimal monotonic extension of it is the set of competitive allocations. [^4] A monotonic extension of the strong core is a superset of the strong core that satisfies Maskin’s Monotonicity, which is necessary for Nash implementation.
@@
Alcalde-Unzu and Molis (2011) and Jaramillo and Manjunath (2012) independently propose the first generalizations of TTC to full preferences that preserve Pareto efficiency and strategy-proofness. Aziz and De Keijzer (2012) study the computational complexity of these algorithms, and Plaxton (2013) proposes another algorithm with enhanced computational efficiency. Saban and Sethuraman (2013) unify these generalizations by defining a family of algorithms and derive sufficient conditions on pointing rules to preserve strategy-proofness. Ahmad (2021) study weak group strategy-proofness of these algorithms. Our generalization of TTC to find elements of our proposed core concepts follow the common formats of these existing algorithms. But because we do not pursue strategy-proofness, our description shows more flexibility in selecting pointing rules.
Sandholtz and Tai (2025) study the special setting of the housing market model we consider in Section 3.1, in which agents agree on their indifferent preferences. They show that TTC with fixed tie-breaking is Pareto efficient and core-selecting. This is implied by our results, since the exclusion core is Pareto efficient and we show that it coincides with the strong core whenever the latter is nonempty and it coincides with the set of all TTC outcomes in that special setting. Compared to arbitrary tie-breaking, the benefit of fixed tie-breaking is that it ensures group strategy-proofness of TTC in that special setting.
@@
There is a literature studying standard core concepts in the housing market model with externalities under strict preferences (Mumcu and Saglam, 2007; Graziano et al., 2020; Hong and Park, 2022; Doğan et al., 2011; Aslan and Lainé, 2020; Klaus and Meo, 2023). It shows that even the weak core can be empty under unrestricted strict preferences. So, this literature typically focuses on restricted preference domains so that the strong core or other solutions can be nonempty.
Finally, following BK, a few papers further study the exclusion core under strict preferences. Balbuzanov and Kotowski (2024) generalize that concept to a model incorporating production. Cheng et al. (2024) develop that concept in a priority-based allocation model. Ishida and Park (2025) study a special case of BK’s original model in which agents are partitioned into groups and each group collectively owns a partitioned group of objects. Zhang (2025) proposes a refinement of the exclusion core in BK’s original model. We are the first to examine the exclusion core under weak preferences.